OpenStudy (anonymous):

Find the transverse axis of 45y^2-320x^2+6=2886

OpenStudy (sidsiddhartha):

\[\frac{ y ^{2} }{ a ^{2} }-\frac{ x ^{2} }{ b ^{2} }=1\] try to write it in this form first

OpenStudy (sidsiddhartha):

any ideas?

OpenStudy (anonymous):

1/64y^2-1/9x^2=1?

OpenStudy (sidsiddhartha):

yup correct

OpenStudy (sidsiddhartha):

now length of transverse axis of a hyperbola like this is=2*a so what's your (a)?

OpenStudy (anonymous):

1/64?

OpenStudy (sidsiddhartha):

 do u get it?